a drive in cashier took $20 in bills to the bank to get change. she asked for twice as many dimes as nickels and three times as many quarters as nickels. Was the teller able to grant her request?
help

Found 2 solutions by bucky, stanbon:
Answer by bucky(2189)

Let N equal the number of nickels.
.
Since the number of dimes is to be twice the number of nickels, the number of dimes is 2*N.
.
And since the number of quarters is to be three times the number of nickels, the number of
quarters is 3*N
.
Since there are N nickels and each nickel is $0.05, the amount of money from nickels will
be 0.05*N.
.
Since there are 2*N dimes and each dime is $0.10, the amount of money from dimes will be
(0.10)*2*N which equals 0.2*N.
.
Finally since there are 3*N quarters and each quarter is $0.25, the amount of money from
quarters will be (0.25)*3*N = 0.75N.
.
Now we need to find the total amount of money from nickels, dimes, and quarters by adding
these three amounts and setting it equal to $20. In equation form this is:
.
0.05*N + 0.2*N + 0.75*N = 20
.
Add the three terms on the left side. They total 1.0*N so the equation becomes:
.
1.0*N = 20
.
Divide both sides by 1.0 and you get N = 20
.
So there are 20 nickels.
.
Therefore the number of dimes is twice the number of nickels, so there are 40 dimes.
.
And the number of quarter is three times the number of nickels, so there are 60 quarters.
.
Now let's check it out. 20 nickels is $1.00. 40 dimes is $4.00. And 60 quarters (at 4 quarters
per dollar) is $15. This total is $1.00 + $4.00 + $15.00 = $20.00. The answer of 20 nickels,
40 dimes, and 60 quarters is correct.
.
Hope this helps you to understand the problem and how you can get the answer.

Answer by stanbon(75887)

a drive in cashier took $20 in bills to the bank to get change. she asked for twice as many dimes as nickels and three times as many quarters as nickels. Was the teller able to grant her request?
-------------
Let # of nickels be "x" ; value is 5x cents
# of dimes = "2x"; value is 10*2x = 20x cents
# of quarters = "3x"; value is 25*3x = 75x cents
----------------------
EQUATION:
value + value + value = 2000 cents
5x + 20x + 75x = 2000
105x = 2000
x = 2000/105 = 19.047...
--------------
x is not a whole number so the teller cannot satisfy the request.
===========
Cheers,
Stan H.